OpenStudy (anonymous):

The difference of two positive numbers is 9. The larger is twice the smaller decreased by 5. Find the two numbers

OpenStudy (anonymous):

This Problem is similar to your previous one, so you can just use the same method that was used before.

OpenStudy (anonymous):

let x and y be these numbers. Stament a) x-y=9 Stat. b) x=2(y-5) using x (b) and subst in (a) 2(y-5)-y=9 2y-10-y=9 y=19 using this value in (a) x-19=9 x=28 Solution x=28; y=19. I think that`s it....